OpenStudy (anonymous):

If 1 of the 100 subjects is randomly selected, find the probability of getting someone who did not have a headache.

OpenStudy (anonymous):

the data: 10mg Atorvastatin Placebo Headache 15 65 No Headache 17 3

OpenStudy (anonymous):

\[\begin{array}{l|c|c|} &10\text{mg Atorvastatin}&\text{Placebo}&\color{red}{\text{Total}}\\ \hline \text{Headache}&15&65&\color{red}{80}\\ \hline \text{No Headache}&17&3&\color{red}{20}\\ \hline \color{blue}{\text{Total}}&\color{blue}{32}&\color{blue}{68}&\color{purple}{100}\\ \hline \end{array}\] According to the table, 80 of the 100 subjects have a headache. Knowing this, what's the probability that a random subject chosen will have a headache?

OpenStudy (anonymous):

I don't know if i'm doing this correctly, but the probability with headache: 80/100 (simplified to 4/5) and with no headache: 20/100 (simplified 1/5).

OpenStudy (anonymous):

Yes, \(\dfrac{80}{100}=\dfrac{4}{5}\).